OpenStudy (anonymous):

What is the sum of a 38–term arithmetic sequence where the first term is 14 and the last term is 154?:)

OpenStudy (asnaseer):

this one requires a slightly different formula than your last question.

OpenStudy (asnaseer):

do you know the formula for the sum of a series?

OpenStudy (anonymous):

I know a=14 and n=38. but whats d?

OpenStudy (ghazi):

s= n/2(A1+An)

OpenStudy (asnaseer):

you don't need to calculate d

OpenStudy (asnaseer):

there are at least two forms for the sum of an arithmetic series...

OpenStudy (asnaseer):

\[S_n=\frac{n}{2}(2a+(n-1)d)\]and:\[S_n=\frac{n}{2}(a+l)\]where "l" is the last term of the series.

OpenStudy (anonymous):

so, x38=38/2(14+154)?

OpenStudy (asnaseer):

yes

OpenStudy (asnaseer):

but not x38 - it is \(S_{38}\)

OpenStudy (anonymous):

so is it 3192?:)

OpenStudy (asnaseer):

if you recall from your previous question, you had:\[x_n=a+(n-1)d\]to get the n'th term

OpenStudy (ghazi):

yes

OpenStudy (ghazi):

3192

OpenStudy (asnaseer):

so if you look at the first equation for the sum of a series we can see:\[S_n=\frac{n}{2}(2a+(n-1)d)\]\[\qquad=\frac{n}{2}(a+a+(n-1)d))\]\[\qquad=\frac{n}{2}(a+x_n)\]

OpenStudy (asnaseer):

that is how the second form of the sum is derived.

OpenStudy (asnaseer):

hope that makes sense?